OpenStudy (anonymous):

simplify the rational expression x-3/x^2-5x+6? help

OpenStudy (compassionate):

Hi, \[\frac{ x - 3 }{ x^2 - 5x + 6 }\] You want to factor the bottom term. \[\frac{ x - 3 }{ (x - 2)(x - 3) }\] The x - 3 will cancel out, leaving: \[\frac{ 1 }{ x-2 }\]

OpenStudy (compassionate):

OpenStudy Ambassador, OpenStudy Team.

OpenStudy (anonymous):

thankyou so much (: